OpenStudy (anonymous):

A new freezer costs $500 plus $0.20 a day to operate. An old freezer cots $20 plus $0.50 a day to operate. After how many days in the cost of operating each freezer equal? Solve by using row reduction

OpenStudy (anonymous):

it can be either 500 days 20 days 820 days 1600 days

OpenStudy (rulnick):

500+0.2x=20+0.5x 480 = 0.3x x=1600

OpenStudy (rulnick):

... after 1600 days each would cost $820.

OpenStudy (anonymous):

thx buddy! u rock

OpenStudy (rulnick):

welcome